OpenStudy (anonymous):

Help... let f(x)=2x and g(x)=5-3x. Find,[a] g(x)-f(x) [b] f(-1) [c] g(0)-f(-1).

Parth (parthkohli):

\( \color{Black}{\Rightarrow g(x) - f(x ) = (5 - 3x) - (2x)}\) \( \color{Black}{\Rightarrow f(-1) = 2(-1)}\) \( \color{Black}{\Rightarrow g(0) - f(-1) = (5 - 3(0)) - 2(-1) }\)

Parth (parthkohli):

The key to solving these problems is substitution. I believe that you could continue all 3 of them.

OpenStudy (anonymous):

thanks, but let me try with your example. i will get back to you.

Parth (parthkohli):

Sure.

OpenStudy (anonymous):

i am clear with that. thanks one more time.